# EAS-SNN: End-to-End Adaptive Sampling and Representation for Event-based Detection with Recurrent Spiking Neural Networks

### v1

@@ Abstract
Event cameras, with their high dynamic range and temporal resolution, are ideally suited for object detection, especially under scenarios with motion blur and challenging lighting conditions. However, while most existing approaches prioritize optimizing spatiotemporal representations with advanced detection backbones and early aggregation functions, the crucial issue of adaptive event sampling remains largely unaddressed. Spiking Neural Networks (SNNs), which operate on an event-driven paradigm through sparse spike communication, emerge as a natural fit for addressing this challenge. In this study, we discover that the neural dynamics of spiking neurons align closely with the behavior of an ideal temporal event sampler. Motivated by this insight, we propose a novel adaptive sampling module that leverages recurrent convolutional SNNs enhanced with temporal memory, facilitating a fully end-to-end learnable framework for event-based detection. Additionally, we introduce Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to regulate potential distribution and address performance degradation encountered in spike-based sampling modules. Through rigorous testing on neuromorphic datasets for event-based detection, our approach demonstrably surpasses existing state-of-the-art spike-based methods, achieving superior performance with significantly fewer parameters and time steps. For instance, our method achieves a 4.4% mAP improvement on the Gen1 dataset, while requiring 38% fewer parameters and three time steps. Moreover, the applicability and effectiveness of our adaptive sampling methodology extend beyond SNNs, as demonstrated through further validation on conventional non-spiking detection models.
@@
However, the inherent asynchronous and sparse characteristics of data captured by event cameras present significant obstacles for traditional dense neural networks, which are primarily designed to process synchronous, uniformly structured data. To reconcile the asynchronous event stream with the dense tensor formats utilized in ANNs, most existing frameworks incorporate a sampling-aggregation mechanism, also referred to as event representation or embedding [30, 18, 3, 19, 38]. This integration effectively alleviates the computational demands engendered by the high temporal resolution of event data. Building upon this foundation, researchers have delved into exploiting the unique spatiotemporal attributes of event data through refining event representations [31, 52, 32, 34, 8] and enhancing the design of detection architectures [41, 46, 32, 53, 49, 19, 22], thereby boosting network efficacy. Additionally, the aggregation module is integrated with downstream feature extraction components into a unified end-to-end training framework [18]. Regarding event-based sampling, several studies have endeavored to enhance the sampling module from a rule-based or search-based perspective [32, 38, 69, 5]. Nonetheless, these initiatives frequently suffer from the lack of high-level semantic information or are encumbered by artificial constraints and suboptimal search efficiency. So far, the quest for a differentiable, adaptive sampling module that facilitates end-to-end optimization within event-based detection frameworks remains unfulfilled.
@@
On the other hand, SNNs, composed of biologically plausible spiking neurons, support asynchronous event-driven computation on neuromorphic architectures[1, 9, 45] and so present a high potential for fast inference and low power consumption. This positions SNNs as inherently compatible with the sparse nature of event data. Moreover, a spectrum of low-level representations [52, 31, 18] potentially allows for their encapsulation within a unified SNN layer [28]. Nevertheless, the challenge of training deep SNNs under long time steps persists, in spite of rapid advancements in spike-based learning algorithm [58, 51, 39, 62, 14, 35, 68, 55]. This makes it difficult to align the discretization window of SNNs with event response time, culminating in a discernible information temporal information loss. Therefore, the current event-based detection models based on SNNs exhibit a notable performance discrepancy when compared to their ANN counterparts.
@@
To address these challenges, this work introduces ARSNN, a learnable adaptive sampling methodology with recurrent convolutional SNNs (Fig. 1a), enabling a comprehensive end-to-end learning paradigm for event representation and feature extraction. This methodology utilizes pixel-wise spiking neurons in the sampling module to capture the dynamic information delivered by events across receptive fields. Upon the information integration surpassing a predefined threshold, the events between the last and the current spike firing times are sampled as shown in Fig. 1b. By optimizing the spike firing time and membrane potential distribution, the sampling module can be trained and coordinated with downstream spike-based detectors, thereby greatly mitigating the temporal alignment loss between the raw event stream and SNNs. Moreover, the proposed SNN-based approach is optimally positioned for deployment on neuromorphic hardware, potentially curtailing overhead from data migration. The main contributions are outlined as follows:
@@
We first propose a learnable adaptive sampling method and realize an end-to-end event-based detection framework via SNNs that seamlessly integrates the processes of sampling, aggregation, and feature extraction.
@@
We develop the techniques of Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to mitigate the performance degradation problem in SNN-based sampling modules with the recurrent convolutional architecture.
@@
Empirical evaluations on neuromorphic datasets (N-Caltech 101 and Gen1) show the proposed method achieves superior performance over the state-of-the-art spike-based methods with less energy consumption compared to ANNs. Furthermore, the effectiveness of our adaptive sampling method is verified across dense detection models.
@@ 2.1 Spiking Neural Networks
Spiking Neural Networks (SNNs) have raised substantial interest due to their asynchronous computation paradigm and promising energy efficiency. However, optimizing SNNs poses challenges, particularly due to the binary nature of spike firing. The advancements in backpropagation techniques for SNNs, including time-based [43, 64, 42, 63, 67] and activation-based methods [58, 51, 62, 20, 7, 35, 55], estimate gradients for firing time and spike generation successfully. Additionally, various spiking neuron models [14, 21, 60, 36, 6, 24] have been proposed to enhance network heterogeneity and performance. To maintain gradient flow in deep SNNs, specialized network modules [65, 11, 14] and architectures [13, 53, 27, 59, 66] have been developed considering temporal dynamics in spiking neurons. These collective efforts are narrowing the performance gap between SNNs and ANNs in various tasks, including object detection [61, 53], point cloud recognition [48], and automatic speech recognition [57].
@@ 2.2 Event-based Detection
In event-based detection, methodologies are bifurcated into two principal categories: dense and sparse representations. Dense representations typically convert event data into frame-based formats, subsequently processed by conventional dense detection algorithms [30, 3, 4, 19, 38]. This approach, while facilitating the utilization of pre-trained models within an event-based context [40], tends to significantly compress the inherent spatiotemporal richness of events, necessitating sophisticated event-to-frame conversion techniques. In pursuit of more effectively extracting the spatiotemporal information from event streams, researchers have integrated the recurrent connections [46, 32, 19] and attention mechanisms [54, 33]. Recent advancements [19, 38, 22] strive to minimize detection latency without substantial performance degradation, through the strategic redesign of network backbones and event aggregation modules, thus enhancing adaptability to scenarios demanding high-speed processing.
Conversely, sparse representations, embodied by Graph Neural Networks (GNNs) and SNNs, offer substantial energy efficiency due to their inherent sparse computational models. GNN models [34, 49] construct dynamic spatio-temporal graphs from event stream, updating node and edge states with incoming events. In parallel, SNN methodologies [8, 61, 53], characterized by the internal temporal dynamics and spike-driven computation, naturally complement the asynchronous nature of event stream. However, given the high temporal resolution of event cameras, event-based sampling is imperative to mitigate computational demands and enhance network generalization capabilities, even for sparse representation.
For event-based sampling, a spectrum of methodologies has emerged to refine rule-based strategies, taking into account the instantaneous rate of event generation [32, 29, 38]. Nevertheless, these strategies routinely encounter challenges associated with hyperparameter sensitivity, as well as the difficulty of maintaining an optimal balance between sparse and ambiguous textures in complex scenarios. Nikola et al. [69] have explored the optimization of sampling windows based on the Gromov-Wasserstein discrepancy from raw event data to their representations. Nonetheless, the technique has yet not fully integrated high-level semantic guidance. Cao et al. [5] have proposed the use of SNNs as a mechanism for global event triggering, advocating for a feedback-update mechanism that employs pseudo labels generated by downstream detectors to guide SNN training. Despite these advancements, the integration of ANNs and SNNs remains disjointed, bridged solely by pseudo labels, which introduces artificial constraints on the optimization landscape and compromises search efficiency. In this work, we investigate an adaptive event sampling strategy leveraging learnable SNNs, marking a step towards achieving comprehensive end-to-end event-based representation learning.
@@ 3.1 Event-based Detection
Nonetheless, given the high temporal resolution of event cameras, it is challenging for event-based object detectors to process and interpret times per second. Moreover, the necessity for detecting moving objects at microsecond intervals is often not critical, as object motion typically occurs at a slower pace than the pixel response rate of event cameras [46]. Consequently, the current body of literature [32, 19, 53, 61] frequently incorporates a sampling operation and an aggregation operation to reduce the volume of sparse and asynchronous event data, thereby facilitating more efficient processing prior to the detection phase:
@@
In this formula, denotes the global time window for sampling, and represents the interval of sampling slices. We designate the collection of events that meet the criteria for fixed window sampling within the -th step as . For the -th step sampling, a typical aggregation function, , based on event count histograms, is defined as follows:
Here, represents the Dirac delta function, which equals 1 only when . By employing this approach, the asynchronous and sparse event data are effectively down-sampled into a compact form that significantly alleviates the computational burden on the detector. However, existing sampling techniques often operate independently of downstream object detectors or multi-scale feature extractors, necessitating extensive hyperparameter tuning and showing limited adaptability.
@@ 3.2 Reccurent Convolutional SNNs for Adaptive Event Sampling
In this study, our goal is to develop a learnable sampling module, denoted as , utilizing SNNs to address the challenge of non-differentiable, end-to-end learning that constrains rule-based sampling methods.
The leaky-integrate-and-fire (LIF) model, popular in neuromorphic computing due to its minimal computational requirements, serves as the foundation for our approach. This model is mathematically expressed as follows:
@@
This leaky-integrate-and-fire dynamic closely aligns with the desired properties of an event sampler. A subset of events should be selectively sampled as the accumulated information from the asynchronous event stream reaches a certain threshold. Consequently, we associate the spike firing in SNNs with event-based sampling, integrating this functionality into an end-to-end optimization framework as shown in Fig. 2.
Firstly, restricted by the poor support for sparse spike-driven computation on GPUs, we perform early aggregation using event count as shown in Eq. 4 with a small time slice and obtain the actual -step inputs for . It is worth noting that in general, the closer the timestamps of events are to the detection target timestamp , the greater the impact on the detection result at . Therefore, in this case, the input closer to in time is preferentially sent to . This operation is also consistent with the temporal information dynamics in the SNNs which has been observed that information becomes highly concentrated in earlier few timesteps [26]. Following early aggregation, a distinct spiking neuron is allocated to each spatial position and polarity , with each neuron exhibiting unique behavior by accumulating its spatially localized input over time.
where denotes a constant decay factor. Our empirical findings suggest that an effective sampler must possess an advanced temporal credit assignment capability to extract informative and discriminative substreams from the inherently noisy and fluctuating spatiotemporal event data. Consequently, we incorporate recurrent synapse connections, as depicted in the bottom left portion of Fig. 2, to enhance the temporal representational capacity of SNNs:
Here, the spike firing at the last time step contributes to the input current and decay factor at the current time step directly through recurrent synapse and . denotes the sigmoid function which normalizes the decay factor . We refer to SNNs after this modification as “RSNN” compared to Eq. 9, which effectively helps improve detection performance. When the spike is emitted with over-threshold membrane potential, the event slice between the last spike firing time and the current spike firing time is sampled, so as to achieve local sampling:
This approach infuses the adaptability into event sampling, driven by the inherent variability in firing times across spiking neurons located at distinct spatial coordinates. Notably, we adopt the hard reset to clear the membrane potential from the last sample. Following the event-based sampling phase, the subsequent challenge is to effectively integrate each sampled set into a coherent embedding. Empirically, we find that decoupling the sampling and the aggregation modules might hinder the training phase due to unstable gradient flow. Hence, in this work, we unify the sampling and the aggregation within a singular SNN. We directly employ the accumulation of membrane potential within the specified sampling window as the aggregation input for the downstream network:
Such a model incorporating the spike-driven sampling mechanism and the synaptic recurrent connection is referred to as the ‘ARSNN’ for the latter discussion.
@@ 3.3 Mitigating Degradation of Spike-based Sampling
In light of the adaptive sampling, it can be seamlessly integrated with the downstream detector for end-to-end training via backpropagation through time (BPTT) [56] with surrogate gradient functions [58, 51, 14]. Nonetheless, this aggregation of Eq. 12 implicitly optimizes the sampling module by relying solely on the aggregation of membrane potentials, which lacks information from the firing threshold for optimizing the spike firing timing. To elucidate, we leverage the gradients from the activation-based technique in SNN training to compute the update of firing time (detailed derivation provided in the Supplementary Material):
This reveals that the adjustment in firing time is solely dependent on the gradient of the membrane potential and the instantaneous change in potential at the firing moment. It disregards the relationship between the firing time update with the firing threshold . To counteract this issue, we introduce Spike-aware Training (SAT), which incorporates the firing spike into the aggregation function:
@@
where is the surrogate gradient function used for circumventing the non-differential problem of the Heaviside function of Eq. 7. Here, represents the adjustment from threshold-based gradient redirection, enhancing the update of firing time especially when near the threshold , thereby achieving a more precise sampling away from critical points.
Moreover, the implementation of a spike-based sampling module introduces the dilemma of handling data after the final spike emission. An intuitive idea to address this involves incorporating the remaining information into the embedding, aiming to reduce information loss. However, this approach leads to a scenario where neurons, even without firing, accumulate substantial information for the downstream process. This inadvertently aggravates sampling degradation by diminishing the incentive for the sampling module to accurately learn spike timing. Hence, we propose the strategy of Residual Potential Dropout (RPD) to exclude all events following the last spike, enhancing the learning of precise firing moments. To illustrate the efficacy of RPD, we show the potential aggregation distribution with and without RPD. The results, as depicted in Fig. 3, generally reveal a more compact feature distribution with RPD. Moreover, employing SAT for gradient guidance (shown in the case of ARSNN+SAT), we observe that sampler training devoid of RPD tends to exhibit mode collapse towards the non-firing spectrum with concentrated negative membrane potentials.

### v2

@@ Abstract
Event cameras, with their high dynamic range and temporal resolution, are ideally suited for object detection in scenarios with motion blur and challenging lighting conditions. However, while most existing approaches prioritize optimizing spatiotemporal representations with advanced detection backbones and early aggregation functions, the crucial issue of adaptive event sampling remains largely unaddressed. Spiking Neural Networks (SNNs), operating on an event-driven paradigm, align closely with the behavior of an ideal temporal event sampler. Motivated by this, we propose a novel adaptive sampling module that leverages recurrent convolutional SNNs enhanced with temporal memory, facilitating a fully end-to-end learnable framework for event-based detection. Additionally, we introduce Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to regulate potential distribution and address performance degradation encountered in spike-based sampling modules. Empirical evaluation on neuromorphic detection datasets demonstrates that our approach outperforms existing state-of-the-art spike-based methods with significantly fewer parameters and time steps. For instance, our method yields a 4.4% mAP improvement on the Gen1 dataset, while requiring 38% fewer parameters and only three time steps. Moreover, the applicability and effectiveness of our adaptive sampling methodology extend beyond SNNs, as demonstrated through further validation on conventional non-spiking models. Code is available at Github.
@@
However, the inherent asynchronous and sparse characteristics of data captured by event cameras present significant challenges for traditional dense neural networks, which are primarily designed to process synchronous, uniformly structured data. To bridge the asynchronous event stream with the dense tensor formats utilized in ANNs, most existing frameworks incorporate a sampling-aggregation mechanism, also referred to as event representation [32, 19, 3, 20, 40] or event embedding. This integration effectively alleviates the computational demands engendered by the high temporal resolution of event data. Building upon this foundation, researchers have delved into exploiting the unique spatiotemporal attributes of event data through refining event representations [33, 56, 34, 36, 8] and enhancing the design of detection architectures [43, 49, 34, 57, 53, 20, 24], thereby boosting network efficacy. Additionally, the aggregation module is integrated with downstream feature extraction components into a unified end-to-end training framework [19]. Regarding event-based sampling, several studies have endeavored to enhance the sampling module from a rule-based or search-based perspective [34, 40, 75, 5]. Nonetheless, these initiatives often suffer from the lack of high-level semantic information or are hindered by artificial constraints and suboptimal search efficiency. So far, the quest for a differentiable, adaptive sampling module that facilitates end-to-end optimization within event-based detection frameworks remains unfulfilled.
@@
On the other hand, SNNs, composed of biologically plausible spiking neurons, support asynchronous event-driven computation on neuromorphic hardware[1, 9, 47] and so present a high potential for fast inference and low power consumption. This inherent compatibility with event data makes SNNs ideally suited for event-based vision tasks. Moreover, a spectrum of low-level representations [56, 33, 19] can be encapsulated within a unified SNN layer [30]. Nevertheless, the challenge of training deep SNNs under long time steps persists, despite rapid advancements in spike-based learning algorithm [63, 55, 67, 15, 37, 74, 59]. This makes it difficult to align the discretization window of SNNs with event response time, culminating in temporal information loss. Consequently, the current event-based detection models based on SNNs exhibit a notable performance gap when compared to their ANN counterparts.
@@
To address these challenges, this work introduces ARSNN, a learnable adaptive sampling methodology with recurrent convolutional SNNs (Fig. 1a), enabling a comprehensive end-to-end learning paradigm for event representation and feature extraction. This methodology utilizes pixel-wise spiking neurons in the sampling module to capture the dynamic information delivered by events across receptive fields. Upon the information integration surpassing a predefined threshold, the events between the last and the current spike firing times are sampled as shown in Fig. 1b. By optimizing the spike firing time and membrane potential distribution, the sampling module can be trained and coordinated with downstream spike-based detectors, thereby greatly mitigating the temporal alignment loss between the raw event stream and SNNs. The main contributions are outlined as follows:
@@
We first propose a learnable adaptive sampling method and implement an end-to-end event-based detection framework via SNNs, integrating the processes of sampling, aggregation, and feature extraction.
@@
We develop the techniques of Residual Potential Dropout (RPD) and Spike-Aware Training (SAT) to mitigate performance degradation in SNN-based sampling modules with the recurrent convolutional architecture.
@@
Empirical evaluations on neuromorphic datasets (N-Caltech 101, Gen1, and 1Mpx) demonstrate that the proposed method outperforms state-of-the-art spike-based methods with less energy consumption compared to ANNs. Furthermore, the effectiveness of our adaptive sampling method is verified across dense detection models.
@@ 2.1 Spiking Neural Networks
Spiking Neural Networks (SNNs) have raised substantial interest due to their asynchronous computation paradigm and promising energy efficiency. However, optimizing SNNs poses challenges due to the binary nature of spike firing. The advancements in backpropagation techniques for SNNs, including time-based [45, 69, 44, 68, 73] and activation-based methods [63, 55, 67, 22, 7, 37, 59], estimate gradients for firing time and spike generation successfully. Additionally, various spiking neuron models [15, 23, 65, 38, 6, 26] have been proposed to enhance network heterogeneity and performance. To maintain gradient flow in deep SNNs, specialized network modules [70, 12, 15] and architectures [14, 57, 29, 64, 71] have been developed considering temporal dynamics in spiking neurons. These collective efforts are narrowing the performance gap between SNNs and ANNs in various tasks, including object detection [60, 57, 66], action recognition [21, 41, 11], point cloud recognition [52], and automatic speech recognition [62].
@@ 2.2 Event-based Detection
In event-based detection, methodologies are bifurcated into two principal categories: dense and sparse representations. Dense representations typically convert event data into frame-based formats, subsequently processed by conventional detection algorithms [32, 3, 4, 20, 40]. This approach, while facilitating the use of pre-trained models within an event-based context [42], compress the inherent spatiotemporal richness of events, necessitating sophisticated event-to-frame conversion techniques. To extract the spatiotemporal information more effectively, researchers have integrated the recurrent connections [49, 34, 20] and attention mechanisms [58, 35]. Recent advancements [20, 40, 24] strive to minimize detection latency without substantial performance degradation, through the strategic redesign of network backbones and event aggregation modules.
Conversely, sparse representations, embodied by Graph Neural Networks (GNNs) and SNNs, offer substantial energy efficiency due to their inherent sparse computational models. GNN models [36, 53] construct dynamic spatio-temporal graphs from event stream, updating node and edge states with incoming events. In parallel, SNN methods [8, 66, 57], characterized by the internal temporal dynamics and spike-driven computation, naturally complement the asynchronous nature of event stream. However, given the high temporal resolution of event cameras, event-based sampling is imperative to mitigate computational demands and enhance network generalization capabilities, even for sparse representation.
For event-based sampling, some works have emerged to refine rule-based strategies by considering the instantaneous rate of event generation [34, 31, 40]. Nevertheless, these strategies frequently encounter challenges associated with hyperparameter sensitivity and the difficulty of maintaining an optimal balance between sparse and ambiguous textures in complex scenarios. Nikola et al. [75] explored the optimization of sampling windows based on the Gromov-Wasserstein discrepancy from raw event data to their representations. Nonetheless, this technique has yet not fully integrated high-level semantic guidance. Cao et al. [5] proposed using SNNs for global event triggering, advocating for a feedback-update mechanism that employs pseudo labels generated by downstream detectors to guide SNN training. Despite these advancements, the integration of ANNs and SNNs remains disjointed, bridged solely by pseudo labels, which introduces artificial constraints on the optimization landscape and compromises search efficiency. In this work, we investigate an adaptive event sampling strategy leveraging learnable SNNs, marking a step towards achieving comprehensive end-to-end event-based representation learning.
@@ 3.1 Event-based Detection
Nonetheless, given the high temporal resolution of event cameras, it is challenging for event-based object detectors to process and interpret times per second. Moreover, the necessity for detecting moving objects at microsecond intervals is often not critical, as object motion typically occurs at a slower pace than the pixel response rate of event cameras [49]. Consequently, most works [34, 20, 57, 66] frequently incorporate a sampling operation and an aggregation operation to reduce the volume of sparse and asynchronous event data, thereby facilitating more efficient processing prior to the detection phase:
@@
where denotes the global window for sampling, and represents the interval of sampling slices. We designate the index set that meets the criteria for fixed window sampling within the -th step as . For the -th step sampling, a typical aggregation function, , based on event count histograms, is defined as follows:
Here, represents the Dirac delta function, which equals 1 only when . By employing this approach, the asynchronous and sparse event data are effectively down-sampled into a compact form that significantly alleviates the computational burden on the detector. However, existing sampling techniques often operate independently of downstream object detectors , necessitating extensive hyperparameter tuning and showing limited adaptability.
@@ 3.2 Reccurent Convolutional SNNs for Adaptive Event Sampling
In this study, our goal is to develop a learnable sampling module (Fig. 2), denoted as , utilizing SNNs to address the challenge of non-differentiable, end-to-end learning that constrains rule-based sampling methods.
The Leaky-Integrate-and-Fire (LIF) model, popular in neuromorphic computing due to its minimal computational requirements. The discrete LIF equations, via the first-order Euler method with a time slice of , are given by:
Here, and denote the membrane potential and input current of the -th neuron in the -th layer, respectively, with as the membrane time constant. Upon receiving spike inputs, a neuron accumulates the input current into its membrane potential , which decays at a constant rate. When the membrane potential exceeds the threshold , a spike is emitted, resetting to . The LIF dynamic closely aligns with the desired properties of event-driven sampling, where events should be selectively sampled as the accumulated information from the asynchronous event stream reaches a threshold. Consequently, we propose to associate spike firing in SNNs with event-based sampling, integrating this functionality into an end-to-end optimization framework.
Firstly, restricted by the poor support for sparse spike-driven computation on GPUs, we perform early aggregation using event count as shown in Eq. 4 with a small time slice and obtain the actual -step inputs for . Following early aggregation, a distinct spiking neuron is allocated to each spatial position and polarity , with each neuron exhibiting unique behavior by accumulating its spatially localized input over time:
where denotes a constant decay factor. Our empirical findings suggest that an effective sampler must possess an advanced temporal credit assignment capability to extract informative and discriminative substreams from the inherently noisy and fluctuating spatiotemporal event data. Therefore, we incorporate recurrent synaptic connections, as depicted in the bottom left portion of Fig. 2, to enhance the temporal representational capacity of SNNs:
@@
This approach optimizes event sampling by converting it into learning proper firing times across spiking neurons at distinct spatial coordinates. After event-based sampling, the remained challenge is to effectively integrate each sampled set into a coherent embedding. Empirically, we find that decoupling the sampling and the aggregation modules can hinder training due to unstable gradient flow. Thus, we unify the sampling and the aggregation within a single SNN, using the potential accumulation within the specified sampling window as the aggregation input for the downstream network:
Notably, we adopt the hard reset to clear the membrane potential from the previous sample. Such a model incorporating the spike-driven sampling mechanism and the synaptic recurrent connection is referred to as the ‘ARSNN’ for the latter discussion.
@@ 3.3 Mitigating Degradation of Spike-based Sampling
In light of the adaptive sampling, it can be integrated with the downstream detector for end-to-end training via backpropagation through time (BPTT) [61] with surrogate gradient functions [63, 55, 15]. However, this aggregation method in Eq. 9 implicitly optimizes the sampling module by relying solely on the aggregation of membrane potentials, which lacks information from the firing threshold for optimizing the spike firing timing. Formally, we leverage the gradients from the activation-based technique in SNN training to estimate the update of firing time (detailed derivation provided in the Supplementary Material):
This reveals that the adjustment in firing time depends solely on the gradient of the membrane potential and the instantaneous change in potential at the firing moment. It disregards the relationship between the firing time update and the firing threshold . To address this issue, we introduce Spike-aware Training (SAT), which incorporates the firing spike into the aggregation function as follows:
@@
where is the surrogate gradient function used to circumvent the non-differential problem of the Heaviside function in Eq. 5. Here, represents the adjustment from threshold-based gradient redirection, enhancing the update of firing time especially when near the threshold , thereby achieving more precise sampling away from critical points. Moreover, the implementation of a spike-based sampling module introduces the dilemma of handling data after the final spike firing. An intuitive idea is to incorporate the remaining information into the embedding to reduce information loss. However, this leads to a scenario where neurons, even without firing, accumulate substantial information for representation. This inadvertently causes sampling degradation by diminishing the incentive for the sampling module to learn accurate spike timing. Hence, we propose the strategy of Residual Potential Dropout (RPD) to exclude all events following the last spike, enhancing the learning of precise firing moments. To illustrate the efficacy of RPD, we show the potential aggregation distribution with and without RPD. The results in Fig. 3 reveal a more compact feature distribution with RPD. Moreover, employing SAT for gradient guidance (as in the case of ARSNN+SAT), we observe that sampler training without RPD tends to exhibit mode collapse towards the non-firing spectrum with concentrated negative membrane potentials.